OpenStudy (anonymous):

x^2/16 + y^2/9 = 1 What type of conic section is this? Is it horizontal or vertical?

OpenStudy (mertsj):

x^1, y^2 or y^1, x^2 = parabola x^2, y^2, coefficients the same= circle x^2, y^2 both positive, coefficients not equal = ellipse x^2, y^2 one positive, one negative = hyperbola